Suppose you have to work 420 minutes at your job tomorrow.
a. What number of minutes corresponds to 1% of the total number of minutes you have to work tomorrow?
b. To find the number in part (a), we can multiply 420 by what number?
c. If you have worked 1% of the total 420 minutes, what percent of the 420 minutes do you have left to work?
d. If you have worked 1% of the total 420 minutes, how many minutes do you have left to work?
e. To find the answer in part (d), we can multiply 420 by what number?

Answer by Theo(13342)

Suppose you have to work 420 minutes at your job tomorrow.

a. What number of minutes corresponds to 1% of the total number of minutes you have to work tomorrow?

1% of 420 = .01 * 420 = 4.2 minutes.


b. To find the number in part (a), we can multiply 420 by what number?

you would multiply 420 by .01 which is 1% / 100.
1% means 1 per hundred = 1/100 = .01

c. If you have worked 1% of the total 420 minutes, what percent of the 420 minutes do you have left to work?

100% - 1% = 99%

you have 99% of the 420 minutes left to work.

420 - 4.2 = 415.8

99% / 100 = .99 * 420 = 415.8

d. If you have worked 1% of the total 420 minutes, how many minutes do you have left to work?

you have 415.8 minutes left to work.


e. To find the answer in part (d), we can multiply 420 by what number?

you would multiply 420 by 99% / 100 = .99.

-----

to find the decimal equivalent of a percent, you divide the percent by 100.

1% is therefore equal to .01.
99% is therefore equal to .99.
100% is therefore equal to 1.

x% of something is therefore equal to x/100 times something.

example:

10% of 100 = .10 * 100 = 10